OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8

OpenStudy (anonymous):

@iambatman @iamboatman

OpenStudy (anonymous):

>_>

OpenStudy (anonymous):

boatman

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

its ur nick naim

OpenStudy (anonymous):

mah naims jeaph

OpenStudy (anonymous):

Slope intercept form is just y = mx+b

OpenStudy (anonymous):

Solve for y in your equation

OpenStudy (anonymous):

K LEMMY TRY GEEZE GIVIN ME ALL THE ANSWERS ALL THE TIME

OpenStudy (anonymous):

GOSH

OpenStudy (anonymous):

wait is this just like that last problem?

OpenStudy (anonymous):

@iambuttman

OpenStudy (anonymous):

No, you just have to do some algebra, isolate y on its own (get y by itself). To do this, you want to eliminate everything on the left and move it on right, so do the opposite operation of what already exists on the left on the left side (so it cancels) and on the right.

OpenStudy (anonymous):

is it y=3/2x÷4?

OpenStudy (anonymous):

y=3/2x÷4

OpenStudy (anonymous):

the ??????'s are division signs

OpenStudy (anonymous):

@ieatbreadman

OpenStudy (anonymous):

What's 8/4?

OpenStudy (anonymous):

o so its y = 3/2 (divided by) 2 ?

OpenStudy (anonymous):

@ilickbatsman

OpenStudy (anonymous):